# Objective-Free Local Learning and Emergent Language Structure in Thinking Machines

### v1

@@ 4.1. The Replay Game (Model Thinking)
  • Parallelism: Each $a_{\alpha}$ contributes independently to inference.
  • Emergent Computation: The $a_{\alpha}$ contribution to inference generalizes beyond local $n$-gram statistics.
@@ 5.2. Key-Value Memory via Fast Embedding Neurons
If $a_{\alpha}$ neurons respond more rapidly than $v_{k}$, we treat them as reaching a fixed point first (Krotov, 2021):
@@
Eqn 23 defines a hierarchical key-value memory, where:
@@
This emergent key-value memory mechanism allows long-term associations to be built from local dynamics alone, forming the basis for structured retrieval and generalization in later inference. Thus the model can generalize beyond local $n$-gram statistics, enabling long-range coherence, partial match inference, and efficient lookup. We refer to this process as recognition [^4] Note that recognition can interfere with tokenization. For instance, the repeated string formed from monk$+$eye, i.e …onkeyemonkeyemon…, uniquely tokenizes back into the two words stated, however, the reader may still recognize monkey or key even though they don’t uniquely tokenize the string, highlighting the competition between fast long-term memory effects with slower tokenizing short-term memory (Eugenio, 2025)..
@@ 5.2.1. Key-Value Memory Example
Consider Eqn 23 with $g^{(n)}=0$, s.t inference is driven purely by $a_{\alpha}$. Suppose we have two learned word embeddings, $\alpha\in\{1,2\}$, corresponding to:
@@
×$\quad\times\left(\tilde{v}_{\bf ca}^{(2)}m^{(3)}_{1,{\bf ca},k}+\tilde{v}_{\bf at}^{(2)}m^{(3)}_{1,{\bf at},k}+\tilde{v}_{\bf cat}^{(3)}m^{(4)}_{1,{\bf cat},k}\right)$
@@
(24) ×$\quad\times\left(\tilde{v}_{\bf do}^{(2)}m^{(3)}_{2,{\bf do},k}+\tilde{v}_{\bf og}^{(2)}m^{(3)}_{2,{\bf og},k}+\tilde{v}_{\bf dog}^{(3)}m^{(4)}_{2,{\bf dog},k}\right).$
@@ 5.3. Compression via Word-Specific Projector Decomposition
(30) $=(\tilde{v}^{(3)}_{\bf cat}+\tilde{v}^{(3)}_{\bf ats})(\tilde{v}^{(3)}_{\bf cat}\tilde{v}^{(2)}_{\bf ca}v_{\bf t}+\tilde{v}^{(3)}_{\bf ats}\tilde{v}^{(2)}_{\bf at}v_{\bf s}):=\hat{m}^{(3)}_{\bf cats}\tilde{P}^{\bf cats}_{3}.$
@@ 5.4. Generalized Compression via Projector Contraction
The concrete example above illustrates how word-specific projector chains can be disentangled by pruning irrelevant branches. We now formalize this procedure using a sequence of algebraic contractions and SVDs (Schollwöck, 2011; Parker et al., 2020; Gourianov et al., 2022), resulting in a compressed representation of the hierarchical projector chain for each word embedding.
@@
This compressed chain retains only the information needed to reconstruct the replay trajectory of word $\alpha$, eliminating all unrelated feature branches. The compression is exact if no finite SVD eigenvalues are discarded. Thus this naturally generalizes to approximate compressions via truncating irrelevant contributions to the SVD spectra. We leave this latter point for future discovery.
@@
This word-specific decomposition achieves two goals:
@@
  • Compression: Optimizes the dimension of projection maps; & can regulate embedding dimension (approx).
@@ 5.5. Super-Hierarchical Key-Value Memory
and apply bounded updates: $Q\leftarrow f_{Q}(\Delta Q)$ for $Q\in\{v_{k},\ a_{\alpha},\ b_{\xi}\}$. We treat $b_{\xi}$ $a_{\alpha}$ as being faster than $v_{k}$ (i.e hierarchy of equilibriation times $\tau_{b}\ll\tau_{a}\ll\tau_{v}$ (Krotov and Hopfield, 2021; Krotov, 2021)). Thus we solve for $b_{\xi}$ and $a_{\alpha}$ first, then substitute into $\Delta v_{k}$.
@@
5.5.1. Plasticity, Emergence, and the role of Space & Time
@@
The emergent key-value memories discussed here are a consequence of the plastic nature of Hebbian learning, combined with the hierarchical tokenizing inference. These new mechanisms are not put in by hand – there is no “key-value memory layer” added to the system – and no retraining of the entire model is necessary. Instead, introducing a single new neuron to the system is sufficient in order for them to arise, demonstrating that they are truly emergent structures. (Nor is explicit data necessary – see Discussion.)

### v2

@@ 5.1.3. Topology before meaning
  • Parallelism: Each $a_{\alpha}$ contributes independently to inference. Allows for partial staged-key checks.
  • Emergent Computation: The $a_{\alpha}$ contribution to inference generalizes beyond local $n$-gram statistics. The token complexity evolves with continuous learning.
@@ 6.2. Key-Value Memory via Fast Embedding Neurons
If $a_{\alpha}$ neurons respond more rapidly than $v_{k}$, we treat them as reaching a fixed point first (DenseHierarchyHopfield_Krotov2021):
@@
Eqn 29 defines a hierarchical key-value memory, where:
@@
This emergent key-value memory mechanism allows long-term associations to be built from local dynamics alone, forming the basis for structured retrieval and generalization in later inference. Thus the model can generalize beyond local $n$-gram statistics, enabling long-range coherence, partial match inference, and efficient lookup. We refer to this process as recognition [^15] Note that recognition can interfere with tokenization. For instance, the repeated string formed from monk++eye, i.e …onkeyemonkeyemon…, uniquely tokenizes back into the two words stated, however, the reader may still recognize monkey or key even though they don’t uniquely tokenize the string, highlighting the competition between fast long-term memory effects with slower tokenizing short-term memory (Eugenio2025_HebbianLanguage)..
@@ 6.2.1. Key-Value Memory Example
Consider Eqn 29 with $g^{(n)}=0$, s.t inference is driven purely by $a_{\alpha}$. Suppose we have two learned word embeddings, $\alpha\in\{1,2\}$, corresponding to:
@@
$\quad\times\left(\tilde{v}_{\bf ca}^{(2)}m^{(3)}_{1,{\bf ca},k}+\tilde{v}_{\bf at}^{(2)}m^{(3)}_{1,{\bf at},k}+\tilde{v}_{\bf cat}^{(3)}m^{(4)}_{1,{\bf cat},k}\right)$
@@
(30) $\quad\times\left(\tilde{v}_{\bf do}^{(2)}m^{(3)}_{2,{\bf do},k}+\tilde{v}_{\bf og}^{(2)}m^{(3)}_{2,{\bf og},k}+\tilde{v}_{\bf dog}^{(3)}m^{(4)}_{2,{\bf dog},k}\right)$.
@@ 6.3. Compression via Word-Specific Projector Decomposition
(36) $(\tilde{v}^{(3)}_{\bf cat}+\tilde{v}^{(3)}_{\bf ats})(\tilde{v}^{(3)}_{\bf cat}\tilde{v}^{(2)}_{\bf ca}v_{\bf t}+\tilde{v}^{(3)}_{\bf ats}\tilde{v}^{(2)}_{\bf at}v_{\bf s}):=\hat{m}^{(3)}_{\bf cats}\tilde{P}^{\bf cats}_{3}.$
@@ 6.4. Generalized Compression via Projector Contraction
The concrete example above illustrates how word-specific projector chains can be disentangled by pruning irrelevant branches. We now formalize this procedure using a sequence of algebraic contractions and SVDs (mps_Schollwock2011; mps_Parker&Cao&Zaletel2020; mps_Gourianov&all2022_turbulence), resulting in a compressed representation of the hierarchical projector chain for each word embedding.
@@
Because each projector has its own word ($\alpha$) & length ($n$) specific indices, it’s more correct to write $\hat{P}_{n,n^{\prime}}^{(\alpha)}$. Each $\hat{m}^{(n)}$-capped chain having their own set of projectors:
$\hat{m}^{(n)}_{\alpha}\hat{P}_{n,n}^{(\alpha)}\hat{P}_{n,n-1}^{(\alpha)}\cdots\hat{P}_{n,2}^{(\alpha)}.$
This compressed chain retains only the information needed to reconstruct the replay trajectory of word $\alpha$, eliminating all unrelated feature branches. The compression is exact if no finite SVD eigenvalues are discarded.
@@
This word-specific decomposition achieves three goals:
@@
  • Compression: Optimizes the dimension of projection maps.
  • Finer Staged-Key Searches: Reduces the overhead of content addressable memory retrieval.
Warning: One drawback of this compression is that the resulting projectors no longer form a gauge-invariant set. Consequently, it’s not possible to regauge between right/left tokenized forms. This happens because the compression preserves only those DAG subpaths related to merging tokens on the right (i.e $\tilde{v}^{(n-1)}v\rightarrow\tilde{v}^{(n)}$).
This is still an exact compression (keep all SVD eigenvalues), because not all subpaths are being used to project up onto a given $m^{(n)}$—squeezing out only the essential part of the dense representation. The key part of the key-value mechanism remains unaffected; only the redundancy related to extra freedom in how the value part decodes its latent information is discarded. The exact symbolic behavior & boundary information remain. And the entire string can still be decoded exactly from only right-token prediction. Thus it is possible to move back-&-forth between an $\alpha$-specific gauge-invariant projector set $\{P_{n}^{(\alpha)}\}$ & the hyper-compressed $\{\hat{P}_{n,n^{\prime}}^{(\alpha)}\}$.
@@ 6.5. Super-Hierarchical Key-Value Memory
and apply bounded updates: $Q\leftarrow f_{Q}(\Delta Q)$ for $Q\in\{v_{k},\ a_{\alpha},\ b_{\xi}\}$. We treat $b_{\xi}$ & $a_{\alpha}$ as being faster than $v_{k}$ (i.e hierarchy of equilibriation times $\tau_{b}\ll\tau_{a}\ll\tau_{v}$ (HopfieldNeurobiology_Krotov&Hopfield2021; DenseHierarchyHopfield_Krotov2021)). Thus we solve for $b_{\xi}$ and $a_{\alpha}$ first, then substitute into $\Delta v_{k}$.
@@
6.5.1. Plasticity & Emergence
@@
The emergent key-value memories discussed here are a consequence of the plastic nature of Hebbian learning, combined with the hierarchical tokenizing inference. These new mechanisms are not put in by hand – there is no “key-value memory layer” added to the system – and no retraining of the entire model is necessary. Nor is explicit structured data required. Instead, introducing a single new neuron to the system is sufficient in order for them to arise, demonstrating that they are truly emergent structures.